# First Place Solution of 2023 Global Artificial Intelligence Technology Innovation Competition Track 1

### v1

@@ 1 Introduction
In recent years, with the advancement of artificial intelligence technology [23, 25, 4, 22], researchers have begun to explore how to use deep learning models [26, 17, 12] to automatically generate accurate diagnostic reports. These models typically generate natural language descriptions based on sequence generation models such as Recurrent Neural Networks (RNN) [27] or Transformers [21] and require a large amount of annotated data, including descriptions of medical imaging diagnoses and corresponding expert-written diagnostic reports.
Due to data desensitization, it is not feasible to directly fine-tune open-source pre-trained models. Therefore, we select Chinese CPT-BASE [19] as our base model and append the desensitized numbers to the vocabulary of the pre-trained model. To reduce the gap between pre-train and downstream tasks, we remove Masked Language Modeling task from CPT-BASE and adopt a span mask [11] strategy to perform Denoising Auto-Encoding pre-training task, with an increasingly larger mask ratio to enlarge the difficulty of the pre-training tasks. This is beneficial for the pre-training tasks to capture deeper levels of textual context information.
@@ 2.1 Text Generation in NLP
Text generation [29, 10, 8, 28, 1] is a widely researched direction in the field of natural language processing. It involves using computer systems to generate text that resembles human language, and it has extensive applications in areas such as machine translation, intelligent customer service, literary creation, and more. In recent years, significant progress has been made in text generation techniques, which mainly include text summarization and text generation. In the field of text summarization, algorithms based on TF-IDF and sequence-to-sequence (Seq2Seq) models from deep learning are commonly used methods. TF-IDF measures the importance of words in documents, while Seq2Seq models consist of an encoder and a decoder for generating summaries. Additionally, the large language model based on the Transformer architecture in natural language generation has also gained increasing attention. It implements text generation through attention mechanism and encoder-decoder structure.
@@ 3.2 Fine-tuning Stage
It is worth noting that with each iteration of retrieval augmentation, a new retrieved Diagnostic report will be added to the end of the sample. After n iterations, at least 0 pseudo Diagnostic reports will be added, but up to n pseudo Diagnostic reports could be appended. While more Diagnostic reports can bring greater diversity of information, it also means that the sample will contain more noisy information. So, how can we add as many pseudo Diagnostic reports as possible while also ensuring that the model is influenced as little as possible by the noise?

### v2

@@ 1 Introduction
In recent years, with the advancement of artificial intelligence technology [24, 29, 4, 23], researchers have begun to explore how to use deep learning models [30, 18, 12] to automatically generate accurate diagnostic reports. These models typically generate natural language descriptions based on sequence generation models such as Recurrent Neural Networks (RNN) [31, 33, 28] or Transformers [22, 14, 27, 26] and require a large amount of annotated data, including descriptions of medical imaging diagnoses and corresponding expert-written diagnostic reports.
Due to data desensitization, it is not feasible to directly fine-tune open-source pre-trained models. Therefore, we select Chinese CPT-BASE [20] as our base model and append the desensitized numbers to the vocabulary of the pre-trained model. To reduce the gap between pre-train and downstream tasks, we remove the Masked Language Modeling task from CPT-BASE and adopt a span mask [11] strategy to perform the Denoising Auto-Encoding pre-training task, with an increasingly larger mask ratio to enlarge the difficulty of the pre-training tasks. This is beneficial for the pre-training tasks to capture deeper levels of textual context information.
@@ 2.1 Text Generation in NLP
Text generation [34, 10, 8, 32, 1] is a widely researched direction in the field of natural language processing. It involves using computer systems to generate text that resembles human language, and it has extensive applications in areas such as machine translation, intelligent customer service, literary creation, and more. In recent years, significant progress has been made in text generation techniques, which mainly include text summarization and text generation. In the field of text summarization, algorithms based on TF-IDF and sequence-to-sequence (Seq2Seq) models from deep learning are commonly used methods. TF-IDF measures the importance of words in documents, while Seq2Seq models consist of an encoder and a decoder for generating summaries. Additionally, the large language model based on the Transformer architecture in natural language generation has also gained increasing attention. It implements text generation through an attention mechanism and encoder-decoder structure.
@@ 3.2 Fine-tuning Stage
It is worth noting that with each iteration of retrieval augmentation, a new retrieved Diagnostic report will be added to the end of the sample. After n iterations, at least 0 pseudo Diagnostic reports will be added, but up to n pseudo Diagnostic reports could be appended. While more Diagnostic reports can bring a greater diversity of information, it also means that the sample will contain more noisy information. So, how can we add as many pseudo Diagnostic reports as possible while also ensuring that the model is influenced as little as possible by the noise?